OpenStudy (anonymous):

solve for r r^2+3r=7

OpenStudy (anonymous):

r^2+3r-7=0 use quadratic formula = -b+/-\[\sqrt{b^2-4ac}\]/ (2a) where a=1 b=3 c=-7